OpenStudy (ilovebmth1234):

HELP!!! QUICK! ILL GIVE A MEDAL a salesperson at roaring waves beach supplies, Carissa receives a monthly base pay plus commission on all that she sells. if she sells $600 worth of merchandise in one month, she is paid $408. if she sells $900 of merchandise in one month she is paid $462. Find Carissa's salary when she sells $3,700 worth of merchandise.

OpenStudy (lastdaywork):

Let b be the base salary c be the fraction representing commission on sales "...if she sells $600 worth of merchandise in one month, she is paid $408..." Corresponding equation b + 600c = 408 "... if she sells $900 of merchandise in one month she is paid $462..." Can you tell me the corresponding equation for this ^^ sentence.

OpenStudy (ilovebmth1234):

b+900c=462

OpenStudy (ilovebmth1234):

is the final equation 0.18x+300??

OpenStudy (ilovebmth1234):

@LastDayWork

OpenStudy (lastdaywork):

We have two equations and two unknowns; b + 600c = 408 b+900c=462 Find b and c; then "Find Carissa's salary when she sells $3,700 worth of merchandise."

OpenStudy (lastdaywork):

BTW what do you mean by "final equation" ??

OpenStudy (ilovebmth1234):

when i had to do these we used 3 equations... idk i forget all this... but B=300 and C= 0.18

OpenStudy (lastdaywork):

Yep, the third equations is b + 3700c = x We need to find x :)

OpenStudy (ilovebmth1234):

x=966

OpenStudy (lastdaywork):

Yep! It's correct :)

OpenStudy (ilovebmth1234):

cool xD thanks i forgot a lil how to do all this

OpenStudy (lastdaywork):

You're welcome :)